∫(x 2 +4x+4)/(x 2 +5x+7)(x+2)½

∫(x2+4x+4)/(x2+5x+7)(x+2)½

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear tejas akole

∫(x2+4x+4)dx/(x2+5x+7)(x+2)½

=∫dx(x+2)2/(x2+5x+7)(x+2)½

let x+2 =t2

dx =2 tdt

=2∫dt(t)4/(t4+t2+1)

=2∫dt-2∫(t2+1)dt/(t4+t2+1)

=2t-2∫(1+1/t2)dt/{(t-1/t)2 +3}

now put (t-1/t) =u

               (1+1/t2)dt=du

=2t-2∫du/(u2+3)   +c

now solve and put tha value ou u and t in terms of x


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best tejas.
 
Regards,
Askiitians Experts
Badiuddin